Use the present value formula to determine the amount to be invested now or the present value when needed. the desired accumulated amount is $3500 after 10 yearsinvested in an account with 4% interest compounded quarterly.
Answer by Boreal(15235)

3500=P{1+(.04/4))^40, dividing the interest now by 4 and raising to the 4*10 power.
3500=P(1.01)^40=1.4889
without rounding, divide 3500 by 1.01^40=$2350.79